I am having trouble with some weird/difficult proportion questions~
6/25 is equal to e/80
e=?

Found 2 solutions by Alan3354, richwmiller:
Answer by Alan3354(69443)

I am having trouble with some weird/difficult proportion questions~
1.6/25 is equal to e/80
Use the equal sign.
6/25 = e/80
Cross multiply:
25e = 480
e = 96/5
------------
Do the others the same way.
email me at Moral Loophole@aol.com for assistance, or to check your work.
----------------------------
2. 3/8 is equal to 50/g
g=?
3.7/9 is equal to j/22.5
This could be problem #3, or it could be 3.7/9 = j/22.5
j=?
4.6/13 is equal to 7.8/m
m=?

Answer by richwmiller(17219)

Some website rules you won't like:
You are limited to one problem per submission. [One problem at a time.]
To be sure we understand each other, one means one not two or three or more.
No more than 4 submissions daily. [4 problems a day]
No similar problems. [Different type problems only.]
None of those is weird nor difficult.
But you should write them in math talk.
6/25=e/80
25e=6*80
e=6*80/25
e=480/25